OpenStudy (yacoub1993):

Use the graph of the function f, plotted with a solid line, to sketch the graph of the given function g.

OpenStudy (yacoub1993):

\[g(x)=\left| x+5 \right|\]

OpenStudy (isaiah.feynman):

Last one

OpenStudy (yacoub1993):

An investment is worth $3518 in 1995. By 2000 it has grown to $5553. Let y be the value of the investment in the year x, where x = 0 represents 1995. Write a linear equation that models the value of the investment in the year x. \[A. y = -407x + 7588\] \[B. y = \frac{ 1 }{ 407} x + 3518\] \[C. y = -407x + 3518\] \[D. y = 407x + 3518\]

OpenStudy (yacoub1993):

is D correct

OpenStudy (isaiah.feynman):

Not sure. Sorry

OpenStudy (yacoub1993):

@Isaiah.Feynman Use the graph of y = f(x) to graph the given function g. g(x) = -2f(x)

OpenStudy (isaiah.feynman):

last one again

OpenStudy (yacoub1993):

@Isaiah.Feynman Use the graph to determine the x- and y-intercepts. A. x-intercept: -3; y-intercepts: 3, 1, 5 B. x-intercepts: -3, 1, -5; y-intercept: -3 C. x-intercepts: 3, 1, 5; y-intercept: -3 D. x-intercept: -3; y-intercepts: -3, 1, -5